OpenStudy (anonymous):

4/5x35=38 use the numbers 4 5 and 35 to write this problem with two different factors that result in he same product of 28

OpenStudy (anonymous):

thx for this

OpenStudy (coolguy34):

no problem

OpenStudy (anonymous):

i ment 28 not 38

OpenStudy (coolguy34):

oh ok

OpenStudy (coolguy34):

this one is a little bit harder

OpenStudy (anonymous):

yep its the make sure u know what ur doin question lol

OpenStudy (coolguy34):

oh ok

OpenStudy (coolguy34):

now this one has me stumped

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

a hint: 4*7 = 28, so how would you combine 35 & 5 to get 7

OpenStudy (anonymous):

beleve it or not im tryin to figure it out to

OpenStudy (coolguy34):

peachpi please answer my question too

OpenStudy (anonymous):

35+5=40 u cant have it equal 7 peachpi

OpenStudy (anonymous):

ok so you know addition won't work. Run through the other operations. Try division

OpenStudy (anonymous):

i suck at division man

OpenStudy (anonymous):

35:5=7?

OpenStudy (anonymous):

i used calculator

OpenStudy (anonymous):

so the answer is 35:5

OpenStudy (anonymous):

+4

OpenStudy (anonymous):

35 ÷ 5 = 7

OpenStudy (anonymous):

then 7 * 4 = 28

OpenStudy (anonymous):

yep but it needs to be 28 so i add 4 rite

OpenStudy (anonymous):

oh i times

OpenStudy (anonymous):

am i rite

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yay thx i medal and fan u thx so much :)

OpenStudy (anonymous):

bye